# InteLiPlan: An Interactive Lightweight LLM-Based Planner for Domestic Robot Autonomy

### v3

@@ IV-B Evaluation of Scalability Across Tasks
In this experiment, we evaluate the scalability of the approach by 101 task instructions from SayCan [6]. To accommodate the task variants, we fine-tuned Mistral with the tasks outlined in Table III, which covers the expected actions required for the dataset solutions.
This set of experiments added a metric of Execution success rate, which is the ratio of the trials where the robot successfully executes the planned actions. Given that SayCan [6] does not have failure recovery ability, we only report the result of cases without failure recovery. Generative result from PaLM 8B [38] is also reported as a lightweight LLM baseline.
@@
Table IV presents the tested result of the frameworks with the dataset from Saycan. As a 540B model, the SayCan result explains its excellent decision-making capability with throughout understanding of the world. Its trained affordance value provides sufficient embodied knowledge, ensuring high success rate in execution. Besides, we note that prompting with PaLM 8B only successfully plans 38% cases. Mistral-Prompt with our modular structure helps the LLM model to gain embodied intelligence, with the planning success rate to 59%. It is observed that Mistral-Prompt’s failure cases come from its lack of sense of the operating environment. For example, some of the results tell the robot to ‘go to store’ to pick something up, despite being told that it is working in a domestic environment. It is proven that with 7B parameters, our InteLiPlan obtains similar results in comparison to the state-of-the-art 504B SayCan model with 83% planning success rate.
@@
IV-C Evaluation of Cross-embodiedment with Real-robot Experiments
We validate the cross-embodiment feature of the system on two different physical robot platforms that are different in both camera setups and the number of degrees of freedom (DoF). In both robot setups, InteLiPlan is processed entirely onboard with a single NVIDIA Geforce RTX 2070, verifying onboard computing capability. We tested a pre-trained InteLiPlan model on both platforms:
@@ IV-C Evaluation of Cross-embodiedment with Real-robot Experiments
  • No-failure task: the robot is asked to complete a task by processing a high-level command. In this task, we guarantee that the object is within the observation and is reachable.
@@
Table V recorded the success rates of InteLiPlan in generating the correct sequences of action upon the corresponding case. The success rates were also affected by the success of vision and feasibility modules, leading to the failed cases. The result demonstrates that with one model, planned action sequences guarantee executability across embodiments. Fig. 1 and Fig. 6 showcase InteLiPlan results in real-world settings, which include tasks with and without failure recovery. The multi-step failure verifies that our method enables autonomy with seamless end-user intervention.
@@
Figure 6: Demonstration of the interactive failure recovery capability on the physical HSR.
@@
HSR 6e-6 5 1.5
Anytree 8e-6 5.5 1.3
The experiment validates the cross-embodiedment feature that requires no extra finetuning, further strengthening that InteLiPlan’s planning capability is universal to different robots, and also different camera views, as opposed to vision-language-action models [39].
@@ V Conclusion
This paper reduces the effort of defining traditional task planning domains in robotics, provides an intuitive, robot-independent data structure and boosts applicability in resource-constrainted applications. Notably, deploying our approach with the lightweight Mistral 7B model achieves both comparable results with the SOTA baseline and –notably to our approach– real-time onboard computing. Our system was validated on both the physical HSR and Anymal D-Unitree Z1 robots, verifying the cross-embodiedment capability across different DoFs and camera views.
Limitations and Future Work - Our framework is structured as a multimodal system with a feasibility check, however, it does not have a motion-level replanning feature. Incorporating low-level reactions together with high-level replanning capabilities in the form of a dual process can potentially improve failure recovery speed. Furthermore, the sampled path formed by the reachability graph can be further developed into a motion planning subpart of our approach, resulting into a full task and motion planning stack.

### v4

@@ IV-A1 Comparison of Different Methods in Feasible Tasks
In our first experiment, we assume successful outcomes in both object detection and robotic motion feasibility by placing the requested object in the camera view and within the reach of the robot. In this set, the planner is solely required to generate the appropriate action sequences based on human inputs. An example of such conversations is provided in Fig. 3a.
Despite explicit instructions to produce solely an action sequence, Mistral-Prompt and Deepseek-Prompt often generate extraneous text. For a fairer comparison, we only extracted the action sequence when determining the success rate, and ignored the lengthy explanation and slightly misused words (e.g., ‘pick up’ vs ‘pick’) for the action API interface. As presented in the w/o failure column of Table II, the results demonstrate a superior performance of our models, achieving 100% success in scenarios with seen commands and objects, and maintaining high effectiveness even with unseen objects or commands. In contrast, Mistral-Prompt and Deepseek-Prompt experienced performance drops in scenarios involving unseen elements, with more sensitivity over unseen textual commands. It is also observed that Deepseek-Prompt behaves better than Mistral-Prompt, while Deepseek-Ours has slightly lower success rates than Mistral-Ours.
IV-A2 Evaluation of Robustness in Failure Explanation
In this test, we evaluate the model’s ability to explicitly describe the cause of planning failures during the recovery process. The failure experiments include all categories outlined in Table I, with examples of ‘planning failure’ detection in Fig. 3b,c. We only include ‘execution failure’ in failure recovery experiments (Section IV-A3), as it is a failure description that is triggered from the controller level.
Figure 4 shows that our method achieved near-perfect success rates in scenarios with seen commands and seen/unseen objects, underscoring its reliability in familiar settings. It is observed that ‘Task confusion’ failure detection obtains lower success rates than multimodal failure detection. This is mainly due to the ‘ambiguous task’ experiments, where the robot fails to distinguish object categories (e.g., fruits, drinks) from the list of detected objects. Our method increases the success rates of failure explanation by more than twice in these cases, proving that fine-tuning triggers the domain-specific understanding of the robots for which factor (e.g., user preference) to consider during planning.
IV-A3 Evaluation of Robustness in Failure Recovery
We further evaluate the framework’s ability to replan based on human instructions. In each trial, a detected failure is inserted into <history>, and the models are expected to generate the appropriate recovery action.
Fig. 5 demonstrates our framework’s performance in replanning in various scenarios. While ‘task confusion’ mainly processes user preference (e.g., user prefers the robot to get a Coke to a 7-Up), ‘planning failure’ and ‘execution failure’ cases assess the ability of mapping language to the list of pre-defined motions (e.g., output ‘open cupboard’ in response to ‘the object is in the cupboard’). The lower success rates in recovering from ‘execution failure’ are due to the longer sequence of recovery than what was provided in the fine-tuning data. In general, prompting methods fails to embody LLM with sufficient robot-domain understanding. Our method significantly increases the performance of LLM models in robotic task failure recovery, indicating robustness in real-world conditions where adaptability and responsiveness to failures are crucial for operational success.
The w/ failure column of Table II reports success rates in scenarios where the planner must succeed in the two-step process of failure detection followed by failure recovery. The results reveal a significantly increased gap between prompt-based models and our approach as opposed to w/o failure cases, underscoring the effectiveness of our method in handling multi-step interaction and ensuring task completion.
@@ IV-B Evaluation of Scalability Across Tasks
We evaluate the scalability of our approach using 101 task instructions from SayCan [3]. To accommodate the task variants, we fine-tuned Mistral with the tasks outlined in Table III, which covers the expected actions required for the dataset solutions. We also introduce the Execution success rate metric, defined as the ratio of the trials where the robot successfully executes the planned actions. Since SayCan [3] does not support failure recovery, we report only results from cases without recovery. Generative results from PaLM 8B [5] are also reported as a lightweight LLM baseline.
@@
Table IV presents the tested results of the frameworks with the dataset from Saycan. As a 540B model, the SayCan result explains its excellent decision-making capability with throughout understanding of the world. Its trained affordance value provides sufficient embodied knowledge, ensuring a high success rate in execution. Besides, we note that prompting with PaLM-8B only successfully plans 38% cases. Mistral-Prompt with our modular structure helps the LLM model to gain embodied intelligence, with the planning success rate to 59%. It is observed that Mistral-Prompt’s failure cases come from its lack of sense of the operating environment. For example, some of the results tell the robot to ‘go to store’ to pick something up, despite being told that it is working in a domestic environment. It is proven that with 7B parameters, our InteLiPlan obtains similar results in comparison to the state-of-the-art 504B SayCan model with an 83% planning success rate.
@@
IV-C Evaluation of Cross-Embodiedment with Real Robots
We validate the cross-embodiment feature of the system on two different physical robot platforms that are different in both camera placements and the number of degrees of freedom (DoF). The skill sets are named identically in the two platforms to allow direct transfer without retuning from the pre-trained model’s output to action execution. In both robot setups, InteLiPlan is processed entirely onboard with a single NVIDIA GeForce RTX 2070, verifying onboard computing capability. We tested a pre-trained model on these platforms:
@@ IV-C Evaluation of Cross-Embodiedment with Real Robots
  • No-failure task: the robot is asked to complete a task by processing a high-level command, and the objects are guaranteed to be within the observation and reachable.
@@
Table V recorded the success rates of InteLiPlan in generating the correct sequences of action upon the corresponding case. The success rates were also affected by the success of vision and feasibility modules, leading to the failed cases. The result demonstrates that, with one model, the planned action sequences generalize effectively across different embodiments without additional fine-tuning. Fig. 1 and Fig. 6 showcase InteLiPlan results in real-world settings, which include tasks with and without failure recovery. The multi-step failure verifies that our method enables autonomy with seamless end-user intervention.
@@
Figure 6: Demonstration of the interactive failure recovery capability on the physical HSR. (a) captures HITL replanning with success in vision but failure in action feasibility, while (b) recovers from being unable to find the targeted object.
@@
HSR 6e-6 $\pm$ 2e-7 5 $\pm$ 0.14 1.5 $\pm$ 0.07
Anytree 8e-6 $\pm$ 2e-7 5.5 $\pm$ 0.12 1.3 $\pm$ 0.08
The experiment validates the cross-embodiedment feature that requires no extra finetuning, further strengthening that InteLiPlan’s planning capability is universal to different robots by not using robot-specific data. As opposed to vision-language-action models [22], the approach allows direct transfer to robot with different camera views.
@@ V Conclusion
This paper reduces the effort of defining traditional task planning domains in robotics, provides an intuitive, robot-independent data structure, and improves applicability in resource-constrained settings. Notably, deploying our approach with the lightweight Mistral 7B model achieves both comparable results with the SOTA baseline and –notably to our approach– real-time onboard computing. InteLiPlan approaches the performance of the 504B PaLM-SayCan system while operating fully onboard with a lightweight 7B model. Our system was validated on the physical HSR and Anymal D-Unitree Z1 robots, verifying the cross-embodiedment capability across different DoFs and camera views.
Limitations and Future Work - Our framework is a multimodal system with a feasibility check, but it currently lacks motion-level replanning feature. Integrating low-level reactive behaviors with high-level replanning in a dual-process manner could improve failure recovery in real-world scenarios. Furthermore, the sampled path from the reachability graph could be extended into a motion planning module, enabling a full task and motion planning stack.